Dividends on Preferred and Common Stock

Yosemite Bike Corp. manufactures mountain bikes and distributes them through retail outlets in California, Oregon, and Washington. Yosemite Bike has declared the following annual dividends over a six-year period ended December 31 of each year: 20Y1, $24,500; 20Y2, $10,250; 20Y3, $125,250; 20Y4, $95,000; 20Y5, $120,000; and 20Y6, $165,000. During the entire period, the outstanding stock of the company was composed of 25,000 shares of cumulative preferred 2% stock, $90 par, and 100,000 shares of common stock, $4 par.

Required:

1. Determine the total dividends and the per-share dividends declared on each class of stock for each of the six years. There were no dividends in arrears at the beginning of 20Y1. Summarize the data in tabular form. If required, round your answers to two decimal places. If the amount is zero, please enter "0".

Preferred Dividends Common Dividends

Year
Total
Dividends

Total

Per Share

Total

Per Share
20Y1 $24,500 $ $ $ $
20Y2 10,250
20Y3 125,250
20Y4 95,000
20Y5 120,000
20Y6 165,000
$ $

2. Determine the average annual dividend per share for each class of stock for the six-year period. Round your answers to two decimal places.

Average annual dividend for preferred stock $ per share
Average annual dividend for common stock $ per share

3. Assuming a market price of $100 for the preferred stock and $5 for the common stock, determine the average annual percentage return on initial shareholders’ investment, based on the average annual dividend per share (A) for preferred stock and (B) for common stock. Round your answers to two decimal places.

Preferred stock %
Common stock %
Answer #1

Solution 1:

Annual Preferred Dividend = 25000 *$90* 2% = $45,000

Preferred Dividends Common Dividends
Year Total Dividends Total Per Share Total Arrears at year end Total Per Share
20Y1 $24,500 $24,500 $0.98 $20,500 $0 $0.00
20Y2 $10,250 $10,250 $0.41 $55,250 $0 $0.00
20Y3 $1,25,250 $1,00,250 $4.01 $0 $25,000 $0.25
20Y4 $95,000 $45,000 $1.80 $0 $50,000 $0.50
20Y5 $1,20,000 $45,000 $1.80 $0 $75,000 $0.75
20Y6 $1,65,000 $45,000 $1.80 $0 $1,20,000 $1.20
Total $2,70,000 $2,70,000

Solution 2:

Average annual dividend for preferred stock = ($270000/ 6 years) / 25000 shares = $1.80 per share

Average annual dividend for Common stock = ($270000/ 6 years) / 100000 shares = $0.45 per share

Solution 3:

Preferred stock = Average annual dividend per share for preferred stock / market price for preferred stock

= $1.80 / $100 = 1.80%

Common stock = Average annual dividend per share for Common stock / market price for Common stock

= $0.45 / $5 = 9.00%